VIBRATIONALLY HOT HCN IN THE LABORATORY AND IRC+10216

\underline{JOHN C. PEARSON}^{a}, SHANSHAN YU, HARSHAL GUPTA and BRIAN J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109.

HCN has historically been used as a tracer of the dense gas in the in interstellar medium. The envelopes of carbon rich asymptotic giant branch stars are generally rich in HCN; however, the large and generally variable infrared flux emitted by the star enormously complicates the interpretation. HCN in IRC+10216 shows an enormous number of masers and lasers pumped by the central star and often enhanced by line overlaps with other abundant molecules such as acetylene in the infrared. A total of seven laser transitions including two previously unreported transitions associated with the 040 − 011 interacting bands have been observed. To understand the astronomical observations a study of the radio frequency discharge plasma of CH_{4} and N_{2} was performed. Rotational transitions of HCN in vibrational states up to 15,000 cm^{−1} have been observed including inverted levels and a number of previously undetected states. The spectra from IRC+10216 and the laboratory are presented.

aA part of this work was performed at the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. Copyright 2010\begin{matrix}c \\ ◯ \end{matrix} California Institute of Technology. All rights reserved.